Brief session 08:35 to 09:30

Connect 09:30am

Open sessions.

Operate.

09:30am
  • BCU47 temperature:
  • start -1.3C.
  • OCAM ok
  • Argos cal unit deployed using the Argos interface.
  • initial hexapod alignment postions:
  • DX Hexapod -2.55 -1.675 -0.561 -280.6 -253.9 Numbers taken from 201902.
  • ResetBCU47
  • The dark seems very bad one side quadrant.
Acquire the PSF.

Found the PSF:
  • -2.55,-1.6,1.45,-150,-200
  • Found not light in the OCAM,. scan some positions upper the hotspot on the CCD47.
  • Alessandro confirm ocam cover on.
10:10 am
  • STOP Luci comm.
  • Alessandro will travel to the mountain.
  • Think to move to SX LBTI.
  • Luci WFS power off.

Startup LBTI both sides.

https://wiki.lbto.org/LBTI/WFSAlignment

SX hexapod:

3.27,2.55,-1,-15,-56

DX hexapod:

-5,-3,1.25,132,-685

DX power up.
  • CCD47 not coming up.
  • Most probably the power supply power connector should be checked.
  • Ther is a lost of vacuum that it affects the alignment, there is spected to have a different positions.
SX power up.
  • OK.
Mountain crew checking the BCU47 power cable connector.

11:54am
  • Done with the BCU47 for LBTI
  • Clear to move Argos swing arm.
Checking CCD47 not life.
  • Power cycle CCD47 controller.
  • Not luck.
  • Power cycle BCU47
  • set 4x4
  • OK
checking Right PSF.

M3 setup for RR mode.

Changed to ARgos cal unit.

PSF in the CCD37.

Right:

Center the PSF using the styages.

Left:

left side ready to close the loop.

Close loop but observed not commands sent.

rght side center pupils.
  • OK.
  • Close loop.
left side close loop.

TN:20221128_192802

DX close loop and some masterdiagnostic glitches.

SX closed.
  • Reclose loop with Dx.
  • Fabio checking the masterdiagnostic DX glitches
  • message_deamon disconnected for the W.
  • W_stop/start_all
12:46PM

ocam not live during a close loop. it went offline when taking a dark. back on line when set the configuration back to 1x1 cropped.

Power cycle bcu2k+ set up 1x1 cropped.

Left side still running.

tail the masterdiagnostic.

close loop
  • OK.
  • Distrubance on 0.5.
  • optimize:20221128_193513
  • TN:20221128_195701
12:57PM

Left->ocam zeros
  • open loop
  • change binning using the AOArbitrator GUI.
  • OK.
  • Optimize:20221128_201043
  • center pupils from FLAO
  • restart the auxloop to have the bad frame monitor.
  • Fabio checking the auxloop, not stopping work.
  • SX tacking TN.
  • _205527,_201609,_201621
Change AO tables for both systems to run 1700Hz.
  • DX ACE: _20221014_500m_V02.txt
  • SX ACE: 500m1700Hz_20221109.txt

change the Argos source:

12.7, 12.7

shutter DX: 108

shutter SX:64

1:28PM
  • New preset DX
  • AO preset with 12.5
  • estimated 12.44
  • close loop.
  • optimize:203213
  • TT 6.92, ho1:0.81, ho2:0.42,

    adsec rip for the second pass of TT

  • adsec skipped.
  • Naopther preset
  • freq:755
  • optmkze:204454
  • TT:4.12,0.81,0.57
  • gopton
  • TN:_204721,_204739,_204754
Preset SX.
  • A0 Preset: 12.6
  • freq:762Hz
  • optimize:203558
  • TT:3.88,0.73,0.63
  • estimated:12.7
  • Stop AO
  • Same preset.
  • 732Hz
  • optmze:204439
  • 3.95,0.99,0.69
  • gopt on
  • TN:_204825,_204838,_204855
1:50PM

Change just above 14.5

Open loop.

DX shutter:93
  • mag ocam 14.4
SX shutter:56
  • mag ocam 14.4
1:54 PM

load disturbance 500 Hz.

Preset AO
  • mag 14.4
  • Bin2.
  • SX close loop,. camera lens x->-100
  • DX close.
    • freq:753
    • optmize:210050
    • TT:2.63,0.62,0.85
    • TN:_210349,210359,
    • SR:30%
  • SX close.
  • freq:758
  • optmize:210521
  • gan:7.65,1,17,0.5
  • SR:50%
  • TN:_210850,_210850,_210919

2:11PM

Open loop.

16.5

shutter SX:47

shutter DX:80

disturbance 200Hz

close loop SX
  • FREq:285
  • bin3
  • mag 16.4
  • optmize:_211446
  • TT4.84,1.97,0.10
  • gopton
  • loop open, argos shutter reset.
  • reclose.
  • optmize:212328
  • 4.19,1.57,0.10
  • freq 313
  • Bin3
  • TN:_212553,_212620,_
  • SR:42
  • TT jitter:50mas
close loop DX.
  • freq:323
  • bin3
  • optimize: 211819
  • gqin:6.91,0.85,0.10
  • TN:_212417,_212434,_212451
  • SR:26
  • TTjitter 30mas
lunch break

2:38PM

Stop WFS.Adsec rest.

Argos cal U parked.

3:50PM

elab intro

-- JuanCarlosGuerra - 28 Nov 2022
Topic revision: r2 - 28 Nov 2022, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback